OpenStudy (anonymous):

Consider this system of linear equations, where k is a real parameter x-2ky=-3 3x+(1-6k)y-(k+2)z=k-9 -2x+(5k-1)y+(1-k)z=2k^2+6 In each case, select all of the following which are true. a) k=-1 b)k=0 c)k=1 d) k<-1 e) k>1 f) k in in the interval (-1,0) g) k is in the interval (0, -1) Any help would be greatly appreciated because I'm completely lost.

OpenStudy (helder_edwin):

what has to happen with the system. does it have to have one or infinite solutions? does it have to be inconsistent?

OpenStudy (anonymous):

After those choices are listed, it asks for values of k such that the system has an infinite solutions, one solution and values for when the system is inconsistent.

OpenStudy (helder_edwin):

do u know determinants?

OpenStudy (anonymous):

No. Its only my second day in the class and we've only covered the basics of linear systems and this question is on the homework to be graded

OpenStudy (helder_edwin):

ok. then you KNOW gaussian reduction of a matrix. right??

OpenStudy (anonymous):

Yes, I know that

OpenStudy (helder_edwin):

great, then tha matrix of your system is \[ \large \left[\begin{array}{ccc|c} 1 & -2k & 0 & -3\\ 3 & 1-6k & -(k+2) & k-9\\ -2 & 5k-1 & 1-k & 2k^2+6 \end{array}\right] \] apply gaussian reduction to it and post the final matrix. don't do gauss-jordan reduction, you r not looking for the solution of the system, just trying to find out if it is or not consistent.

OpenStudy (helder_edwin):

it should take u no more than 5-10 minutes.

OpenStudy (anonymous):

I'm not sure that I did this right but this is what i got: \[\left[\begin{matrix}1 & -2k & 0 &3 \\ 0 & 1 & -k+2 & k\end{matrix}\right]\] \[\left[\begin{matrix}0 & 9k-1& 1-k &2k ^{2} \end{matrix}\right]\] (Thes last row is part of the matrix. I couldn't figure out how to type it into one.

OpenStudy (helder_edwin):

no way.

OpenStudy (helder_edwin):

i got this \[ \large \left[\begin{array}{ccc|c} 1 & -2k & 0 & -3\\ 0 & 1 & -k-2 & k\\ 0 & 0 & -(1-k)^2 & k^2-k \end{array}\right] \] check your work (honestly) and tell me if you get the same.

OpenStudy (helder_edwin):

if it helps you, the operations i used are: (-3)*first row + second row 2*first row + third row (1-k)*second row + third row

OpenStudy (anonymous):

I'm going over the work and I realized I forgot to write the negative sign for the negative 3 in the first row. I made a similar mistake in the second row where I forgot to keep the parentheses around k+2 so my signs were off. The last row is giving me trouble still. Can you explain why you used (1-k)*second row + third row? I feel like I'm missing something obvious.

OpenStudy (anonymous):

Never mind, Now I understand why I have to use (1-k). Now I just find values for k that which make the system have an infinite number of solutions as well as inconsistent and a unique solution right?

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

the worst that can happen is to lose one of the pivots. for which value(s) of k do u lose the thirs pivot?

OpenStudy (helder_edwin):

sorry i made a mistake the last rows should be \[ \large `[0\quad0\quad-(1-k^2)\quad|\quad k^2-k] \] check this.

OpenStudy (helder_edwin):

have to go for a while. if u still need help, i'll be back in 15 minutes.

OpenStudy (anonymous):

I get that for the last row as well. I saw you mentioned pivots, my professor has never even mentioned that word before... thank you so much for your help so far!

OpenStudy (helder_edwin):

i'm back

OpenStudy (helder_edwin):

a "pivot" is the entry in the matrix u use to do the reduction. for instance u used the 1 in the first row to introduce zeros below it, this 1 is your first pivot.

OpenStudy (anonymous):

Ok, I get what that means. But how do you know which values lose their pivots?

OpenStudy (helder_edwin):

ahh. the last pivot which is \(-(1-k^2)\) could turn zero if \(k=\pm1\) agree??

OpenStudy (anonymous):

Yes, I agree there. So does that mean there is infinitely many solutions for that value?

OpenStudy (helder_edwin):

wait. now look at the independent term of the third row: \(k^2-k\) if you put \[ \large k^2-k=0\qquad\Rightarrow\qquad k=0\text{ or }k=1 \] agree??

OpenStudy (anonymous):

Yes, I follow you

OpenStudy (helder_edwin):

look at the coincidences: for the pivot u got \(k=\pm1\) and for the independent term u got \(k=0\vee k=1\) If k=1 then u lose the pivot and the independent term. so u lose the whole third row, and now u have more variables than equations. how many solutions do u have in this case??

OpenStudy (anonymous):

I think you have no solutions because when you have more variables than equations you cannot have a unique solution or an infinite number of them?

OpenStudy (helder_edwin):

infinite solutions.

OpenStudy (helder_edwin):

now if k=-1 you lose the third pivot but u do NOT lose the independent term. so your third row would be like this \[ \large (0\quad 0\quad0\quad|\quad2) \] how many solutions do u have in this case???

OpenStudy (anonymous):

You have no solutions because \[0\neq2\]

OpenStudy (helder_edwin):

yeah!! great.

OpenStudy (helder_edwin):

now finally. if \(k\neq\pm1\) then u don't lose the pivot. this means u have three non-zero pivots in a \(3\times3\) system. how many solutions do u have in this case??

OpenStudy (anonymous):

That means there is one unique solution

OpenStudy (helder_edwin):

yes. fair lady. and we r done!

OpenStudy (anonymous):

ok! so just to make sure I got everything straight, when k is -1 or 0, there is no solution when k is 1, there is an infinite number of solution when k is anything other than 1 or -1, there is one solution. Right? And on the interval between (-1,0), there is infinite solutions?

OpenStudy (helder_edwin):

when k=1, infinite solutions. when k=-1, no solutions. when k is anything other than 1 or -1, one solution. regarding the interval, u have to decide according to the three choices above.

OpenStudy (anonymous):

All right. I think I got it now. Thank you so much for all your help. You are a way better teacher than my professor. Thank you!!

OpenStudy (helder_edwin):

u r welcome.